OpenStudy (jasmine8181):

A concert charges $24.50 per daily adult ticket and $13.75 per child’s ticket. A group of 90 people paid $1560.00 for tickets. Write a system of equations that could be used to find x, the number of adult tickets purchased, and y, the number of children's tickets purchased? PLEASE SHOW WORK

OpenStudy (rcp031):

Let x be the number of adult tickets purchased. Let y be the number of children's tickets purchased. Know let us form equations out of given data, A group of 90 people mean, x+y= 90 A group of 90 people paid $1560.00 for tickets mean 24.50x+13.75y=1560.00 Solve for x and y, x+y= 90 24.50x+13.75y=1560.00

likeabossssssss (likeabossssssss):

welcome to os and plz close this :)